4. Given the following mutually exclusive events U, D, & R:

U = robot moves up, D = robot moves down, R = robot moves right, and event L= robot moves off-board to the left

If P (U ) = .29 , P(L | U) = .25, P(D) = .31, P(L | D) = .12 , P(R) = .40, P (L | R )= .38  

Use the total law of probability to find the probability this robot moves left off the board in the next line of code.
Answer #1

P(Left) = P(L | U) * P(U) + P(L | D) * P(D) + P(L | R) * P(R) = 0.25*0.29 + 0.12*0.31 + 0.38*0.40 = 0.2617